# nonblind {nonblind}

### v1

@@ 3 Making Sense of Recommendations with Potential Outcomes
Having set up a principal–agent problem for thinking about the design of recommendation algorithms, we now use the potential-outcomes framework from causal inference to discuss how the agent’s decision depends on the recommendation algorithm. Viewing the agent’s decision in this way suggests that the optimization of recommendations shares common features with the analysis of instrumental variables in inference problems. We strengthen this connection by applying a monotonicity assumption similar to that of Imbens and Angrist (1994), which allows us to represent the remaining potential outcomes in terms of the active decision the agent makes when no recommendation is given and whether the agent complies with the recommendation. This leads to a decomposition of the objective of the recommendation algorithm, in terms of compliance, into the performance of the recommended actions and the quality of active decisions.
@@ 3.1 Recommendations Induce Potential Outcomes
We now interpret the agent’s decision in terms of the potential-outcomes framework of Rubin (1974). Specifically, we write $D_{U}(r;f^{\textnormal{rec}})$ for the potential decisions that the agent would take when receiving the recommendation $r$ from the recommendation algorithm $f^{\textnormal{rec}}$. This notation makes explicit that potential outcomes vary with the private information $U$, and are specific to a recommendation rule $f^{\textnormal{rec}}$. The realized decisions are then obtained as $D=D_{U}(R;f^{\textnormal{rec}})$, where $R=f^{\textnormal{rec}}(X)$.
@@
In our hiring example, the algorithm’s recommendations impact the company’s utility only through managers’ decisions. Our framework, therefore, parallels instrumental variables that affect outcomes only through their impact on a treatment variable. Rather than directly optimizing for the recommendation, the company must understand how the managers’ decision-response $D_{U}(r;f^{\textnormal{rec}})$ varies across $r\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$. Since decisions are themselves binary and the recommendations take three values, we can enumerate these potential outcomes and categorize them into four response types based on how the recommendation impacts the assisted managers’ decisions. These response types are similar in nature to the types developed by Angrist, Imbens, and Rubin (1996) for the analysis of instrumental variables.
@@
Table 1 enumerates the potential decisions associated with a manager’s response to the recommendations and provides such a categorization. If the manager maintains the same decision regardless of the recommendation, they Ignore the recommendation. If they adopt the recommendation regardless of whether the algorithm recommends not hiring (N) or hiring ( H) the applicant, then they Comply with the recommendation. If, instead, the manager always takes the decision that the algorithm does not recommend, then they Defy the recommendation. If none of these cases hold, then it must be the case that the manager will Change their decision in response to the algorithm sending a recommendation, taking one decision when the algorithm sends no recommendation and taking the opposite decision if either recommendation is sent. While we maintain all these response types for now, we will later simplify the analysis by ruling out the Defy and Change types, analogously to the monotonicity assumption of Angrist, Imbens, and Rubin (1996).
@@ 3.2 Objective of Recommendation Algorithms
Having introduced a framework to capture potential decisions in response to recommendations, we now consider the optimal recommendation algorithm as the solution to an optimization problem that minimizes the loss achieved by the potential outcomes the algorithm induces. We show the optimal recommendation’s objective can be separated into a triage term that assumes that recommendations are implemented directly as well as an objective that captures the agent’s endogenous response to the recommendation. We compare this optimal recommendation algorithm to solutions that instead directly target optimal decisions, and provide an intuition for why that optimal decision rule will rarely be the optimal recommendation rule.
In our hiring example, the optimal recommendation algorithm for the company to implement is the mapping from applicant worker data to recommendations that minimizes the expected loss of the decisions taken by the manager,
@@
$f^{\textnormal{rec}*}=\min_{f:X\mapsto R}\mathbb{E}\left[\ell(Y,D)\right]\text{ where }R=f(X),D=D_{U}(R;f).$ (2)
@@
This optimization problem is quite difficult to solve, given the uncertainty about how managers will respond to recommendations. As a result, companies often implement algorithms that instead recommend the optimal decision based on the information available to the algorithm,
$f^{\textnormal{dec}*}=\min_{f:X\mapsto R}\mathbb{E}\left[\ell(Y,R)\right]\text{ where }R=f(X),$ (3)
implicitly assuming that the recommendation $R=f(X)$ gets implemented directly ($D=R$). An extension suggested in previous work by Raghu et al. (2019) is to triage instances into those where algorithms outperform humans and those where humans outperform algorithms. If we write $D_{0}=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}})$ where $f^{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\equiv{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$ for the decision taken by the manager in absence of a recommendation algorithm, and we assume that the manager takes this decision when not given a recommendation ($f(X)={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$) and otherwise follows the algorithm, then this objective and its solution correspond in our framework to
@@
$f^{\textnormal{tri}*}=\min_{f:X\mapsto R}\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D_{0})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]\text{ where }R=f(X).$ (4)
This objective again implicitly assumes that recommendations are followed perfectly when given and that manager decisions are not affected by the recommendation algorithm when no recommendation is given ($R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$). In practice, the decision for $R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$ may itself be different from the decision $D_{0}$ without a recommendation algorithm, and we may specifically hope it to improve since additional information and effort may be available for those cases (Raghu et al., 2019; Athey, Bryan, and Gans, 2020). Note also that the case of recommending an optimal decision is a special case where the algorithm always makes a recommendation ($R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$), with which the agent complies.
We now compare these different objectives and solutions, and show that the optimal recommendation ($f^{\textnormal{rec}*}$) can be thought of as an augmentation of the optimal triage ($f^{\textnormal{tri}*}$) that also incorporates the compliance behavior of the manager. Specifically, the loss function that defines the optimal recommendation can be decomposed into the loss function of the triage problem (which we call the triage effect, TE) and the response effect (RE) from managers deviating from the triage decision,
@@
$f^{\textnormal{rec}*}=\min_{f:X\mapsto R}$ $\overbrace{\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D_{0})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]}^{\textnormal{{TE} -- Triage Effect}}$ (5)
@@
$+\underbrace{\mathbb{E}\left[\left(\ell(Y,D)-\ell(Y,R)\right)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\left(\ell(Y,D)-\ell(Y,D_{0})\right)\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]}_{\textnormal{{RE} -- Response Effect}}$
@@
In addition to the triage effect from (4), this decomposition has a second part that represents deviations from recommendations and unassisted decisions. It can represent both improvements (in cases where the human decision-maker uses their private information to override a suboptimal recommendation or adjusts their active decision based on their knowledge of the recommendation) as well as mistakes (when the manager goes against a correct hiring recommendation or takes an active decision that is worse than in the case without any recommendations).
The two components also highlight the challenge of learning good recommendations from data. The triage effect (TE) is straightforward to estimate, provided the company has access to historical data through which they can learn the joint distribution of covariates $X$, ability $Y$, and unassisted manager decisions $D_{0}$. On the other hand, the response effect depends on the potential outcomes associated with the managers’ decisions, which complicates its evaluation for any $f^{\textnormal{rec}}$ under which decision data has not already been collected. Understanding $(D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}};f^{\textnormal{rec}}),D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}}),D_{U}(\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}};f^{\textnormal{rec}}))$ may require more than simple experimental evaluation based on local changes in recommendations. These potential decisions may vary arbitrarily with changes to $f^{\textnormal{rec}}$. Solving for the optimal recommendation then requires understanding this distribution across potentially implementable recommendation algorithms $\mathcal{F}\subseteq\{f;f:X\mapsto R\}$. This makes designing an optimal recommendation algorithm an intractable task without additional assumptions (even when we just want to robustly outperform the unassisted decisions $D_{0}$).
@@
Figure 1: Representation of potential algorithms $\mathcal{F}$ that could be implemented in terms of the triage effect TE and the response effect RE. The optimal triage $f^{\textnormal{tri}*}$ minimizes the triage effect over $\mathcal{F}$, while the optimal recommendation rule $f^{\textnormal{rec}*}$ minimizes the sum $TE+RE$. The optimal decision rule $f^{\textnormal{dec}*}$ need not lie on the Pareto frontier that minimizes these two losses, highlighted in red.
The relationship between algorithms that solve for the different objectives above is illustrated in Figure 1. Plotting the set of possible algorithms ($\mathcal{F}$) according to their triage effect (TE) and response effect (RE) clarifies, in particular, the relationship between optimal triage and optimal recommendation. The optimal triage ($f^{\textnormal{tri}*}$) minimizes the triage effect without considering the response effect related to each algorithm. The optimal recommendation ($f^{\textnormal{rec}*}$) potentially incurs additional loss through the triage effect in order to reduce the loss incurred due to the response effect. In practice, this means the optimal recommendation may be sub-optimal if applied directly, instead relying on the manager’s ability to strategically override decisions inconsistent with their own private information ($U$) and preferences. However, both $f^{\textnormal{rec}*}$ and $f^{\textnormal{tri}*}$ lie on the Pareto frontier which simultaneously minimizes the triage and response effects. This does not have to be true of the optimal algorithmic decision rule $f^{\textnormal{dec}*}$, which generally incurs a higher triage effect TE than the optimal triage solution $f^{\textnormal{tri}*}$. This is because the triage solution $f^{\textnormal{tri}*}$ comes from minimizing expected triage loss (4) subject to a less strict constraint than the decision solution $f^{\textnormal{dec}*}$. Meanwhile, the response effect can be better or worse when applying the $f^{\textnormal{dec}*}$ or $f^{\textnormal{tri}*}$ rules.

### v2

@@ 3 Making Sense of Recommendations with Potential Outcomes
Having set up a principal–agent problem for thinking about the design of recommendation algorithms, we now use the potential-outcomes framework from causal inference to discuss how the agent’s decision depends on the recommendation algorithm. Viewing the agent’s decision in this way suggests that the descriptive and prescriptive analysis of recommendations shares common features with the investigation of instrumental variables in inference problems. We strengthen this connection by applying a monotonicity assumption similar to that of Imbens and Angrist (1994), which allows us to represent the remaining potential outcomes in terms of the active decision the agent makes when no recommendation is given and whether the agent complies with the recommendation. This leads to a decomposition of the recommendation algorithm’s objective based on the agent’s compliance choice. The objective captures the performance of the recommended action if the agent complies and the performance of the agent’s active decision otherwise.
@@ 3.1 Recommendations Induce Potential Outcomes
We now interpret the agent’s decision in terms of the potential-outcomes framework of Rubin (1974). Specifically, we write $D_{U}(r;f^{\textnormal{rec}})$ for the “potential decisions” that the agent would take when receiving the recommendation $r$ from the recommendation algorithm $f^{\textnormal{rec}}$ when the agent’s information is $U$. Thus, the agent’s decisions can be seen as the random variable $D=D_{U}(R;f^{\textnormal{rec}})$, where $R=f^{\textnormal{rec}}(X)$. This expanded framework is represented visually in Figure 2, which expands Figure 1 by including potential outcomes.
The potential decisions $D_{U}(r;f^{\textnormal{rec}})$ depend not only on possible recommendation values $r$ of $R$, they also depend on the whole mapping $f^{\textnormal{rec}}$ from characteristics $X$ to recommendations $R=f^{\textnormal{rec}}(X)$. This is because how a specific recommendation is interpreted depends on how recommendations are given overall. For example, if an algorithm recommends hiring for everybody, then it may change individual decisions differently from it recommending to hire only a few applicants. This is related to the concept of interference in causal inference, where potential outcomes may be functions not just of their own treatment assignment, but also of those of others. In Figure 2, the line between the recommendation algorithm and the potential decision clarifies that potential decisions may depend on the overall recommendation algorithm itself. This dependence highlights the challenge in estimating good recommendation algorithms under minimal assumptions or using limited data (which we return to in Section 5). Effective ways of forming recommendation algorithms will, therefore, rely on assumptions that limit the complexity of potential decisions, including their dependence on the overall recommendation algorithm.111 Without additional assumptions, this interference challenge implies, in our case, that the potential decision function $D_{U}(\cdot;f^{\textnormal{rec}})$ for a specific instance $(X,U)$ is only ever evaluated at its implied recommendation $R=f^{\textnormal{rec}}(X)$. However, by assuming that the potential decision function $D_{U}(\cdot;f^{\textnormal{rec}})$ only depends on $U$, this notation also captures variation in recommendations given for instances that are otherwise equivalent from the perspective of the human decision-maker. That is, for a given instance $U$, they capture how decisions vary across possible realizations of $f^{\textnormal{rec}}(X)$ stemming from variation in $X$. This formalization, therefore, implies that two instances that present with the same $U,R$ values are not distinguishable to the decision-maker, leading to the same decision even if their $X$ varies. (Here, any noise in decisions would be captured through $U$.)
$X$$U$$D_{U}(\cdot;f^{\textnormal{rec}})$$D$$R$$Y$$\ell(Y,D)$$f^{\textnormal{rec}}$
Figure 2: This diagram expands Figure 1 by adding potential decisions $D_{U}(\cdot;f^{\textnormal{rec}})$ explicitly, which determine decisions by $D=D_{U}(R;f^{\textnormal{rec}})$. The line between the recommendation algorithm and potential decisions points to the challenge that the latter may be affected by the overall design of the algorithm beyond the instance-specific recommendation.
@@
Our framework parallels instrumental variables that affect outcomes only through their impact on a treatment variable. In our hiring example, the algorithm’s recommendations (corresponding to the instrument) impact the company’s utility (playing the role of the outcome) only through managers’ decisions (treatment). Rather than directly optimizing for the recommendation, the company must understand how the managers’ decisions $D$ vary in response to receiving different recommendations from different algorithms. Specifically, as the manager’s information $U$, the recommendation $R$, and the algorithm $f^{\textnormal{rec}}$ are sufficient information to determine $D$, and we control the recommendation algorithm $f^{\textnormal{rec}}$ (and through it $R$), we write the manager’s decision as the potential outcome $D_{U}(r;f^{\textnormal{rec}})$ for each of the possible values $r\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$ of the recommendation. Since decisions are themselves binary and the recommendations take three values, we can enumerate these potential outcomes into eight triplets which we categorize into four response types based on how the realized recommendation impacts the assisted managers’ decisions. These response types are similar in nature to the types developed by Angrist, Imbens, and Rubin (1996) for the analysis of instrumental variables.
@@
Table 1 enumerates the potential outcomes associated with a manager’s decision response to recommendations and provides the aforementioned categorization. If the manager maintains the same decision regardless of the recommendation, they Ignore the recommendation (similar to always-takers and never-takers in causal inference). If they adopt the recommendation regardless of whether the algorithm recommends not hiring (N) or hiring ( H) the applicant, then they Comply with the recommendation. If, instead, the manager always takes the decision that the algorithm does not recommend, then they Defy the recommendation. If none of these cases hold, then it must be the case that the manager will Change their decision in response to the algorithm sending a recommendation, taking one decision when the algorithm sends no recommendation and taking the opposite decision if either recommendation is sent. While we maintain all these response types for now, we will later simplify the analysis by ruling out the Defy and Change types, analogous to the monotonicity assumption of Angrist, Imbens, and Rubin (1996).
@@ 3.2 Objective of Recommendation Algorithms
Having introduced a framework to capture potential decisions in response to recommendations, we now consider the optimal recommendation algorithm as the solution to an optimization problem that minimizes the loss achieved by the decisions the algorithm induces. We show that the recommendation’s objective can be separated into a triage term that assumes that recommendations are implemented directly and a term that captures the agent’s endogenous response to the recommendation. We compare this optimal recommendation algorithm to solutions that instead directly target optimal decisions, and provide an intuition for why that optimal decision rule will rarely be the optimal recommendation rule.
In general, the optimal recommendation algorithm is the one that minimizes the expected loss of the decision taken by the human decision-maker after seeing the recommendation,
@@
$f^{\textnormal{rec}*}=\underset{f:X\mapsto R}{\mathrm{argmin}}\ \mathbb{E}\left[\ell(Y,D)\right]\text{ where }R=f(X),D=D_{U}(R;f).$ (2)
@@
In our hiring example, the optimal recommendation algorithm for the company to implement is the mapping from applicant worker data to recommendations that leads to good hiring decisions by the manager. As this example illustrates, this optimization problem is generally quite difficult to solve, given the uncertainty about how managers will respond to recommendations. As a result, companies often implement algorithms that instead recommend the optimal decision based on the information available to the algorithm,
@@
$f^{\textnormal{dec}*}=\underset{f:X\mapsto R}{\mathrm{argmin}}\ \mathbb{E}\left[\ell(Y,R)\right]\text{ where }R=f(X),R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},$ (3)
implicitly assuming that the recommendation $R=f(X)$ gets implemented directly ($D=R$) and ignoring the algorithm’s ability to not send a recommendation ($R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$). An extension suggested in previous work by Raghu et al. (2019) is to triage instances into those where algorithms outperform humans and those where humans outperform algorithms. We write $D^{0}_{U}=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}})$ where $f^{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\equiv{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$ for the decision taken by the manager in the absence of a recommendation algorithm. If we assume that the manager takes this decision when not given a recommendation ($f(X)={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$) and that otherwise the recommendation is implemented directly, then this objective and its solution correspond in our framework to
@@
$f^{\textnormal{tri}*}=\underset{f:X\mapsto R}{\mathrm{argmin}}\ \mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D^{0}_{U})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]\text{ where }R=f(X).$ (4)
In practice, recommendations may not be followed perfectly. In addition, even the decision for $R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$ may itself be different from the decision $D^{0}_{U}$ without a recommendation algorithm, and we may specifically hope it to improve since additional information and effort may be available for those cases (Raghu et al., 2019; Athey, Bryan, and Gans, 2020).
We now compare these different objectives and solutions, and show that the optimal recommendation ($f^{\textnormal{rec}*}$) differs from triage ($f^{\textnormal{tri}*}$) in that it also incorporates the compliance behavior of the manager and considers how active decisions may differ from unassisted decisions. Specifically, the loss function that defines the optimal recommendation can be decomposed into the loss function of the triage problem (which we call the triage effect, TE) and the response effect (RE) from managers deviating from the triage decision,
@@
$f^{\textnormal{rec}*}=\underset{f:X\mapsto R}{\mathrm{argmin}}$ $\overbrace{\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D^{0}_{U})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]}^{\textnormal{{TE} -- Triage Effect}}$ (5)
$+\underbrace{\mathbb{E}\left[\left(\ell(Y,D)-\ell(Y,R)\right)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\left(\ell(Y,D)-\ell(Y,D^{0}_{U})\right)\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]}_{\textnormal{{RE} -- Response Effect}}$
@@
In addition to the triage effect from (4), this decomposition has a second part representing deviations from recommendations and unassisted decisions. It can represent both improvements (in cases where the human decision-maker uses their private information to override a suboptimal recommendation or adjusts their active decision based on the knowledge they receive through the recommendation algorithm) as well as mistakes (when the manager goes against a correct hiring recommendation or takes an active decision that is worse than in the case without any recommendations).
The two components also highlight the challenge of learning good recommendations from data. The triage effect (TE) is straightforward to estimate, provided the company has access to historical data through which they can learn the joint distribution of characteristics $X$, ability $Y$, and unassisted manager decisions $D^{0}_{U}$. On the other hand, the response effect depends on the potential outcomes associated with the managers’ decisions, which complicates its evaluation for any $f^{\textnormal{rec}}$ under which decision data has not already been collected. Understanding $D_{U}(r;f^{\textnormal{rec}})|X,Y$ may require more than simple experimental evaluation based on local changes in recommendations. These decision distributions may vary arbitrarily with changes to $f^{\textnormal{rec}}$. Solving for the optimal recommendation then requires understanding this distribution across a set of reasonable recommendation algorithms $\mathcal{F}\subseteq\{f;f:X\mapsto R\}$. This makes designing an optimal recommendation algorithm an intractable task without additional assumptions (even when we just want to robustly outperform the unassisted decisions $D^{0}_{U}$).
@@
Figure 3: Representation of a given set of candidate recommendation algorithms $\mathcal{F}$ that could be implemented in terms of the triage effect TE and the response effect RE. The optimal triage $f^{\textnormal{tri}*}$ minimizes the triage effect over $\mathcal{F}$, while the optimal recommendation rule $f^{\textnormal{rec}*}$ minimizes the sum $TE+RE$. The optimal decision rule $f^{\textnormal{dec}*}$ need not lie on the Pareto frontier that minimizes these two losses, highlighted in red.
The relationship between algorithms that solve for the different objectives above is illustrated in Figure 3. Plotting a set of candidate algorithms ($\mathcal{F}$) according to their triage effect (TE) and response effect (RE) clarifies, in particular, the relationship between optimal triage and optimal recommendation. The optimal triage ($f^{\textnormal{tri}*}$) minimizes the triage effect without considering the response effect related to each algorithm. The optimal recommendation ($f^{\textnormal{rec}*}$) trades off additional loss through the triage effect with loss incurred due to the response effect. In practice, this means the optimal recommendation may be sub-optimal if applied directly and instead relies on the manager strategically overriding recommendations to make better decisions. However, both $f^{\textnormal{rec}*}$ and $f^{\textnormal{tri}*}$ lie on the Pareto frontier that simultaneously minimizes the triage and response effects. This does not have to be true of the optimal algorithmic decision rule $f^{\textnormal{dec}*}$, which generally incurs a higher triage effect TE than the optimal triage solution $f^{\textnormal{tri}*}$. This is because the triage solution $f^{\textnormal{tri}*}$ comes from minimizing expected triage loss (4) subject to a less strict constraint than the decision solution $f^{\textnormal{dec}*}$. Meanwhile, the difference in the response effect between the $f^{\textnormal{dec}*}$ and $f^{\textnormal{tri}*}$ rules may be positive or negative.